OpenStudy (anonymous):

What is the product of the square root of 2. (6bthe square root of 16. - 2b)?

OpenStudy (anonymous):

\[\sqrt{2}(6b \sqrt{16}-2b)\]

OpenStudy (anonymous):

Please simplify if possible.

OpenStudy (tkhunny):

\(\sqrt{16}\) -- You had better tell me if that can be simplified.

OpenStudy (anonymous):

4

OpenStudy (tkhunny):

Got it. Substitute that and tell me what's next.

OpenStudy (anonymous):

substitute it as b?

OpenStudy (tkhunny):

Why would you do that? We have done nothing with 'b'. Rather than write \(\sqrt{16}\), write 4.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

\[\sqrt{2}(6b \sqrt{4}-2b)\]

OpenStudy (tkhunny):

Why is it \(\sqrt{4}\)? \(\sqrt{16} = 4 \ne \sqrt{4}\) Write it again, with just '4'.

OpenStudy (anonymous):

\[\sqrt{2}(6b+4-2b)\]

OpenStudy (tkhunny):

Why did it turn into "+4". It was multiplied by 6b a moment ago. Rewrite it again without adding anything. Just substitute. Write it EXACTLY as it was originally - with ONE difference. Don't write \(\sqrt{16}\). Write 4 instead.

OpenStudy (anonymous):

\[\sqrt{2}(10b -2b)\]

OpenStudy (tkhunny):

Where did the 6 go? That was NOT an EXACT reproduction. Make an EXACT reproduction. Don't make ANY other change. What does \(b\sqrt{16}\) mean? Some number b, multiplied by the number \(\sqrt{16}\). If you make a direct substitution, you should get \(b4\), meaning some number b, multiplied by 4. Don't try to do anything with it. JUST SUBSTITUTE. \(\sqrt{16} = 4\)

OpenStudy (anonymous):

ok ...

OpenStudy (anonymous):

im just confused

OpenStudy (tkhunny):

No, you must learn to substitute. \(6b\sqrt{16} = 6b4\) I agree that's a little awkward-looking. Maybe it would be better if we inserted the multiplication symbols. \(6\cdot b\cdot\sqrt{16} = 6\cdot b\cdot 4\) Does that make sense?

OpenStudy (anonymous):

Yes

OpenStudy (tkhunny):

Excellent. Now, write the whole expression with that substitution. ONLY that substitution.

OpenStudy (anonymous):

\[\sqrt{2}(6b4-2b)\]

OpenStudy (tkhunny):

Got it. Can we now simplify 6*b*4? Maybe 6*4*b = 24b?

OpenStudy (anonymous):

\[\sqrt{2}(24b-2b)\]

OpenStudy (tkhunny):

Can we now simplify what is inside the parentheses?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

What say you?

OpenStudy (anonymous):

do i subtract 24b and 2b?

OpenStudy (tkhunny):

Is that what the expression indicates?

OpenStudy (tkhunny):

Can you subtract them? Are they the same sort of thing?

OpenStudy (anonymous):

i think so their like terms

OpenStudy (tkhunny):

Very good. Subtract away!

OpenStudy (anonymous):

\[\sqrt{2}(22b)\]

OpenStudy (tkhunny):

Are we done or is there more to do.

OpenStudy (anonymous):

We multiply

OpenStudy (tkhunny):

We are simplifying. That is all we are doing. Do you see any multiplication in there that will SIMPLIFY the expression?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

Really? I don't. What are you planning?

OpenStudy (anonymous):

no there is none

OpenStudy (anonymous):

the question what is the product so i thought you had to multiply

OpenStudy (tkhunny):

It is multiplication, but the task is to simplify. When operations fail to simplify, quit. \(\sqrt{2}(22b) = 22\sqrt{2}\cdot b = 22b\sqrt{2}\) They are all equivalent and they are all about the same complexity. I think we're done.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Thank you so much for your help!

OpenStudy (anonymous):

:)

OpenStudy (tkhunny):

No worries. It's what we do here. Thank you for hanging in there and getting all the way through it!

OpenStudy (anonymous):

Your welcome . I understand so much better now.